# SalUn: Empowering Machine Unlearning via Gradient-based Weight Saliency in Both Image Classification and Generation

### 01 Mar 2024 v4

@@ 1 Introduction
Machine unlearning (MU) is the problem of efficiently and effectively mitigating the influence of particular data points on a pre-trained model (Shaik et al., 2023). It emerged in response to data protection regulations like ‘the right to be forgotten’ (Hoofnagle et al., 2019). However, its scope and significance rapidly expand to tackle many trustworthy machine learning (ML) challenges in computer vision (CV). These challenges include the defense against backdoor poisoning attacks (Liu et al., 2022a), the enhancement of model fairness (Oesterling et al., 2023), the refinement of pre-training methods to augment transfer learning capabilities (Jain et al., 2023; Jia et al., 2023), and the prevention of text-to-image generative models from generating sensitive, harmful, or illegal image content when exposed to inappropriate prompts (Gandikota et al., 2023).
@@ 4 Challenges in Current Machine Unlearning Methods
Figure 2: The instability limitations of MU methods on CIFAR-10. (a) Sensitivity of unlearning accuracy gaps with respect to Retrain (measured by $|\text{Method}-\text{{\text{Retrain}}}|$’) as a function of forgetting data amount. Four MU methods (FT, RL, IU, $\ell_{1}$-sparse) are included. (b) Box plots illustrating unlearning accuracy using Retrain, IU, and the proposed weight saliency-integrated IU across various hyperparameter choices. The box size represents the variance of UA against hyperparameter values.
@@
The instability limitation. In evaluating the performance of MU methods, previous research has often assumed a fixed number of forgetting data, such as data points within an entire class or a fixed ratio of the training set. There has been limited evaluation exploring how the unlearning performance is affected by varying quantities of forgetting data. In Fig. 2(a), we investigate the unlearning performance gap relative to the gold standard Retrain, measured in terms of UA (unlearning accuracy), as a function of the quantity of forgetting data points. Note that a smaller performance gap is desirable. As we can see, the unlearning effectiveness of MU methods (①-⑤) observed at a 10% forgetting data quantity does not necessarily hold when the forgetting data quantity is increased to 50%. Similarly, the instability can also be observed in other performance metrics and will show in experiment results later.
@@ 5 SalUn: Weight Saliency Is Possibly All You Need for MU
$\mathbf{m}_{\mathrm{S}}=\mathbf{1}\left(\left|\nabla_{{\bm{\theta}}}\ell_{\mathrm{f}}({\bm{\theta}};\mathcal{D}_{\mathrm{f}})\left.\right|_{{\bm{\theta}}={{\bm{\theta}}_{\mathrm{o}}}}\right|\geq\gamma\right),$ (3)
where $\mathbf{1}(\mathbf{g}\geq\gamma)$ is an element-wise indicator function which yields a value of $1$ for the $i$-th element if $g_{i}\geq\gamma$ and 0 otherwise, $|\cdot|$ is an element-wise absolute value operation, and $\gamma>0$ is a hard threshold. In practice, we have observed that setting $\gamma$ to the median of the gradient vector $\nabla_{{\bm{\theta}}}\ell_{\mathrm{f}}({\bm{\theta}};\mathcal{D}_{\mathrm{f}})\left.\right|_{{\bm{\theta}}={{\bm{\theta}}_{\mathrm{o}}}}$ is a sufficiently effective choice. Based on (3), we express the unlearning model ${\bm{\theta}}_{\mathrm{u}}$} as
${{\bm{\theta}}_{\mathrm{u}}}=\underbrace{\mathbf{m}_{\mathrm{S}}\odot{\bm{\theta}}}_{\text{salient weights}}+\underbrace{(\mathbf{1}-\mathbf{m}_{\mathrm{S}})\odot{{\bm{\theta}}_{\mathrm{o}}}}_{\text{original weights}},$ (4)
@@
$\displaystyle\operatorname*{\text{minimize}}_{{\bm{\theta}}}\penalty 10000\ L_{\text{{SalUn}}}^{(1)}({\bm{\theta}}_{\mathrm{u}})\mathrel{\mathop{:}}=\mathbb{E}_{(\mathbf{x},y)\sim\mathcal{D}_{\mathrm{f}},y^{\prime}\neq y}\left[\ell_{\mathrm{CE}}({\bm{\theta}}_{\mathrm{u}};\mathbf{x},y^{\prime})\right],$ (6)
where $y^{\prime}$ is the random label of the image $\mathbf{x}$ different from $y$, and ${\bm{\theta}}_{\mathrm{u}}$ has been defined in (4). Additionally, to achieve a balance between unlearning on forgetting data points and preserving the model’s generalization ability for non-forgetting data points, we usually fine-tune the original model ${\bm{\theta}}_{\mathrm{o}}$} using (6) for a small number of epochs, e.g., 10 epochs in the classification task.
@@
$\displaystyle\operatorname*{\text{minimize}}_{{\bm{\theta}}}\penalty 10000\ L_{\text{{SalUn}}}^{(2)}({\bm{\theta}}_{\mathrm{u}})\mathrel{\mathop{:}}=\mathbb{E}_{(\mathbf{x},c)\sim\mathcal{D}_{\mathrm{f}},t,\epsilon\sim\mathcal{N}(0,1),c^{\prime}\neq c}\left[\|\epsilon_{{{\bm{\theta}}_{\mathrm{u}}}}(\mathbf{x}_{t}|c^{\prime})-\epsilon_{{{\bm{\theta}}_{\mathrm{u}}}}(\mathbf{x}_{t}|c)\|_{2}^{2}\right]+\alpha\ell_{\mathrm{MSE}}({{\bm{\theta}}_{\mathrm{u}}};\mathcal{D}_{\mathrm{r}}),$ (7)
where $c^{\prime}\neq c$ indicates that the concept $c^{\prime}$ is different from $c$ , ${\bm{\theta}}_{\mathrm{u}}$} is the saliency-based unlearned model given by (4), $\alpha>0$ is a regularization parameter to place an optimization tradeoff between the RL-based unlearning loss over the forgetting dataset $\mathcal{D}_{\mathrm{f}}$ and the diffusion training loss $\ell_{\mathrm{MSE}}({{\bm{\theta}}_{\mathrm{u}}};\mathcal{D}_{\mathrm{r}})$ on the non-forgetting dataset $\mathcal{D}_{\mathrm{r}}$ (to preserve image generation quality). Similar to (6), SalUn begins with the pre-trained model ${\bm{\theta}}_{\mathrm{o}}$} and follows the optimization in (7) to accomplish unlearning.
Extension on ‘soft-thresholding’ SalUn. The implementation of SalUn relies on a pre-selected hard threshold to determine the weight saliency map $\mathbf{m}_{\mathrm{S}}$ in (3). While this hard-thresholding approach performs well, we can also develop an alternative implementation of SalUn that employs soft thresholding and may allow for a more flexible saliency map determination; See Appendix A for more algorithmic details. Yet, in practice, the soft-thresholding variant does not outperform the hard-thresholding version. Thus, we will focus on the hard-thresholding implementation by default.
@@ 6.1 Experiment setups
Data, models, and unlearning setups. In image classification tasks, we focus on random data forgetting and evaluate its performance on the data model (CIFAR-10, ResNet-18) (He et al., 2016; Krizhevsky et al., 2009). Additionally, we extend our evaluation to CIFAR-100 (Krizhevsky et al., 2009), SVHN (Netzer et al., 2011), Tiny ImageNet (Le & Yang, 2015) datasets, and VGG-16 (Simonyan & Zisserman, 2014), Swin-T (Liu et al., 2021) architectures. We also consider the class-wise forgetting setup in image classification. See Appendix B.2 for these additional experiments.
In image generation tasks, we focus on two unlearning scenarios: class-wise forgetting using DDPM with classifier-free guidance (referred to as DDPM) (Ho & Salimans, 2022), and concept-wise forgetting using LDM-based stable diffusion (SD) (Rombach et al., 2022). The class-wise forgetting aims to prevent DDPM from generating images belonging to a specified object class, achieved by using the class name as the diffusion guidance. DDPM-based unlearning experiments will be conducted on the CIFAR-10 dataset. In addition, class-wise forgetting is also considered for SD on the Imagenette dataset (Howard & Gugger, 2020) to prevent image generation from a specified text prompt, which is given by ‘an image of [class name]’. For CIFAR-10, we utilize DDPM for sampling with 1000 time steps. As for Imagenette, we employ SD for sampling with 100 time steps. Unless otherwise specified, the sparsity of weight saliency is set at 50%. Furthermore, we will consider concept-wise forgetting in SD to avoid the generation of NSFW (not safe for work) content, where the concept is given by, for example, a nudity-related prompt (like ‘Shirtless Putin at pride’). We refer readers to Appendix B.1 for other MU training details.
@@ 6.2 Experiment results
Tench 100.00 2.54 90.40 1.13 63.26 1.63
English Springer 98.00 4.68 93.80 1.03 43.12 1.75
Cassette Player 100.00 3.11 99.40 1.56 83.12 2.07
Chain Saw 100.00 2.43 94.00 2.48 53.43 2.09
Church 100.00 0.67 45.80 2.09 20.93 1.45
French Horn 100.00 2.27 99.60 1.54 36.87 1.33
Garbage Truck 100.00 1.60 89.60 7.40 52.50 2.57
Gas Pump 100.00 2.57 91.40 2.80 57.65 2.28
Golf Ball 96.00 2.03 94.20 1.12 11.87 1.39
Parachute 100.00 3.08 76.20 1.42 46.40 1.32
Average 99.40 2.49 87.44 2.26 46.92 1.79
@@
Furthermore, we justify the effectiveness of the proposed weight saliency map ($\mathbf{m}_{\mathrm{S}}$) in (3) in Appendix B.2 through the integration of $\mathbf{m}_{\mathrm{S}}$ into classical MU methods (FT, RL, GA, and IU).
@@
Extended from Fig. 4, we quantify the unlearning performance of Retrain, ESD, and SalUn using the two metrics introduced earlier, FID and UA, in Appendix B.3. With a comparable UA performance, it is worth highlighting that SalUn significantly outperforms ESD in FID. ESD seems to exhibit instability when forgetting and learning low-quality images like CIFAR-10. We also investigate the selection of the sparsity ratio for weight saliency in Appendix B.1.
@@ 8 Acknowledgement
C. Fan, J. Liu, and S. Liu were supported by the Cisco Research Faculty Award and the National Science Foundation (NSF) Robust Intelligence (RI) Core Program Award IIS-2207052.
@@
Appendix A Soft-thresholding SalUn

### 04 Apr 2024 v5

@@ 1 Introduction
Machine unlearning (MU) is the task of efficiently and effectively mitigating the influence of particular data points on a pre-trained model (Shaik et al., 2023). It emerged in response to data protection regulations like ‘the right to be forgotten’ (Hoofnagle et al., 2019). However, its scope and significance rapidly expand to tackle many trustworthy machine learning (ML) challenges in computer vision (CV). These challenges include the defense against backdoor poisoning attacks (Liu et al., 2022a), the enhancement of model fairness (Oesterling et al., 2023), the refinement of pre-training methods to augment transfer learning capabilities (Jain et al., 2023; Jia et al., 2023), and the prevention of text-to-image generative models from generating sensitive, harmful, or illegal image content when exposed to inappropriate prompts (Gandikota et al., 2023).
@@ 4 Challenges in Current Machine Unlearning Methods
Figure 2: The instability limitations of MU methods on CIFAR-10. (a) Sensitivity of performance gaps with respect to Retrain (measured by $|\text{Method}-\text{{\text{Retrain}}}|$’) as a function of forgetting data amount. Five MU methods (FT, RL, GA, IU, $\ell_{1}$-sparse) are included. (b) Box plots illustrating unlearning accuracy using Retrain, IU, and the proposed weight saliency-integrated IU across various hyperparameter choices. The box size represents the variance of UA against hyperparameter values.
@@
The instability limitation. In evaluating the performance of MU methods, previous research has often assumed a fixed number of forgetting data, such as data points within an entire class or a fixed ratio of the training set. There has been limited evaluation exploring how the unlearning performance is affected by varying quantities of forgetting data. In Fig. 2(a), we investigate the unlearning performance gap relative to the gold standard Retrain, measured in terms of the average over all the metrics (including UA, RA, TA, and MIA), as a function of the quantity of forgetting data points. Note that a smaller performance gap is desirable. As we can see, the unlearning effectiveness of MU methods (①-⑤) observed at a 10% forgetting data quantity does not necessarily hold when the forgetting data quantity is increased to 50%. Similarly, the instability can also be observed in other performance metrics and will show in experiment results later.
@@ 5 SalUn: Weight Saliency Is Possibly All You Need for MU
$\mathbf{m}_{\mathrm{S}}=\mathds{1}\left(\left|\nabla_{{\bm{\theta}}}\ell_{\mathrm{f}}({\bm{\theta}};\mathcal{D}_{\mathrm{f}})\left.\right|_{{\bm{\theta}}={{\bm{\theta}}_{\mathrm{o}}}}\right|\geq\gamma\right),$ (3)
where $\mathds{1}(\mathbf{g}\geq\gamma)$ is an element-wise indicator function which yields a value of $1$ for the $i$-th element if $g_{i}\geq\gamma$ and 0 otherwise, $|\cdot|$ is an element-wise absolute value operation, and $\gamma>0$ is a hard threshold. In practice, we have observed that setting $\gamma$ to the median of the gradient vector $\nabla_{{\bm{\theta}}}\ell_{\mathrm{f}}({\bm{\theta}};\mathcal{D}_{\mathrm{f}})\left.\right|_{{\bm{\theta}}={{\bm{\theta}}_{\mathrm{o}}}}$ is a sufficiently effective choice. Based on (3), we explicitly express the unlearning model ${\bm{\theta}}_{\mathrm{u}}$} as
${{\bm{\theta}}_{\mathrm{u}}}=\underbrace{\mathbf{m}_{\mathrm{S}}\odot(\Delta{\bm{\theta}}+{{\bm{\theta}}_{\mathrm{o}}})}_{\text{salient weights}}+\underbrace{(\mathbf{1}-\mathbf{m}_{\mathrm{S}})\odot{{\bm{\theta}}_{\mathrm{o}}}}_{\text{original weights}},$ (4)
@@
$\displaystyle\operatorname*{\text{minimize}}_{\Delta{\bm{\theta}}}\penalty 10000\ L_{\text{{SalUn}}}^{(1)}({\bm{\theta}}_{\mathrm{u}})\mathrel{\mathop{:}}=\mathbb{E}_{(\mathbf{x},y)\sim\mathcal{D}_{\mathrm{f}},y^{\prime}\neq y}\left[\ell_{\mathrm{CE}}({\bm{\theta}}_{\mathrm{u}};\mathbf{x},y^{\prime})\right]+\alpha\mathbb{E}_{(\mathbf{x},y)\sim\mathcal{D}_{\mathrm{r}}}\left[\ell_{\mathrm{CE}}({\bm{\theta}}_{\mathrm{u}};\mathbf{x},y)\right],$ (6)
where $y^{\prime}$ is the random label of the image $\mathbf{x}$ different from $y$, and ${\bm{\theta}}_{\mathrm{u}}$ has been defined in (4). Additionally, to achieve a balance between unlearning on forgetting data points and preserving the model’s generalization ability for non-forgetting data points, the regularization term on $\mathcal{D}_{\mathrm{r}}$ preserved, with $\alpha>0$ as a regularization parameter.
@@
$\displaystyle\operatorname*{\text{minimize}}_{\Delta{\bm{\theta}}}\penalty 10000\ L_{\text{{SalUn}}}^{(2)}({\bm{\theta}}_{\mathrm{u}})\mathrel{\mathop{:}}=\mathbb{E}_{(\mathbf{x},c)\sim\mathcal{D}_{\mathrm{f}},t,\epsilon\sim\mathcal{N}(0,1),c^{\prime}\neq c}\left[\|\epsilon_{{{\bm{\theta}}_{\mathrm{u}}}}(\mathbf{x}_{t}|c^{\prime})-\epsilon_{{{\bm{\theta}}_{\mathrm{u}}}}(\mathbf{x}_{t}|c)\|_{2}^{2}\right]+\beta\ell_{\mathrm{MSE}}({{\bm{\theta}}_{\mathrm{u}}};\mathcal{D}_{\mathrm{r}}),$ (7)
where $c^{\prime}\neq c$ indicates that the concept $c^{\prime}$ is different from $c$ , ${\bm{\theta}}_{\mathrm{u}}$} is the saliency-based unlearned model given by (4), $\beta>0$ is a regularization parameter similar to $\alpha$ in (6), to place an optimization tradeoff between the RL-based unlearning loss over the forgetting dataset $\mathcal{D}_{\mathrm{f}}$ and the diffusion training loss $\ell_{\mathrm{MSE}}({{\bm{\theta}}_{\mathrm{u}}};\mathcal{D}_{\mathrm{r}})$ on the non-forgetting dataset $\mathcal{D}_{\mathrm{r}}$ (to preserve image generation quality). Similar to (6), SalUn begins with the pre-trained model ${\bm{\theta}}_{\mathrm{o}}$} and follows the optimization in (7) to accomplish unlearning. See Appendix A for the algorithmic implementations.
Extension on ‘soft-thresholding’ SalUn. The implementation of SalUn relies on a pre-selected hard threshold to determine the weight saliency map $\mathbf{m}_{\mathrm{S}}$ in (3). While this hard-thresholding approach performs well, we can also develop an alternative implementation of SalUn that employs soft thresholding and may allow for a more flexible saliency map determination; See Appendix B for more algorithmic details. Yet, in practice, the soft-thresholding variant does not outperform the hard-thresholding version. Thus, we will focus on the hard-thresholding implementation by default.
@@ 6.1 Experiment setups
Data, models, and unlearning setups. In image classification tasks, we focus on random data forgetting and evaluate its performance on the data model (CIFAR-10, ResNet-18) (He et al., 2016; Krizhevsky et al., 2009). Additionally, we extend our evaluation to CIFAR-100 (Krizhevsky et al., 2009), SVHN (Netzer et al., 2011), Tiny ImageNet (Le & Yang, 2015) datasets, and VGG-16 (Simonyan & Zisserman, 2014), Swin-T (Liu et al., 2021) architectures. We also consider the class-wise forgetting setup in image classification. See Appendix C.2 for these additional experiments.
In image generation tasks, we focus on two unlearning scenarios: class-wise forgetting using DDPM with classifier-free guidance (referred to as DDPM) (Ho & Salimans, 2022), and concept-wise forgetting using LDM-based stable diffusion (SD) (Rombach et al., 2022). The class-wise forgetting aims to prevent DDPM from generating images belonging to a specified object class, achieved by using the class name as the diffusion guidance. DDPM-based unlearning experiments will be conducted on the CIFAR-10 dataset. In addition, class-wise forgetting is also considered for SD on the Imagenette dataset (Howard & Gugger, 2020) to prevent image generation from a specified text prompt, which is given by ‘an image of [class name]’. For CIFAR-10, we utilize DDPM for sampling with 1000 time steps. As for Imagenette, we employ SD for sampling with 100 time steps. Unless otherwise specified, the sparsity of weight saliency is set at 50%. Furthermore, we will consider concept-wise forgetting in SD to avoid the generation of NSFW (not safe for work) content, where the concept is given by, for example, a nudity-related prompt (like ‘Shirtless Putin at pride’). We refer readers to Appendix C.1 for other MU training details.
@@ 6.2 Experiment results
Tench 100.00 2.53 99.40 1.22 42.40 1.63
English Springer 100.00 0.79 100.00 1.02 27.20 1.75
Cassette Player 99.80 0.91 100.00 1.84 93.80 0.80
Chain Saw 100.00 1.58 96.80 1.48 48.40 0.94
Church 99.60 0.90 98.60 1.91 23.80 1.32
French Horn 100.00 0.94 99.80 1.08 45.00 0.99
Garbage Truck 100.00 0.91 100.00 2.71 41.40 0.92
Gas Pump 100.00 1.05 100.00 1.99 53.60 1.30
Golf Ball 98.80 1.45 99.60 0.80 15.40 1.05
Parachute 100.00 1.16 99.80 0.91 34.40 2.33
Average 99.82 1.22 99.40 1.49 42.54 1.30
@@
Furthermore, we justify the effectiveness of the proposed weight saliency map ($\mathbf{m}_{\mathrm{S}}$) in (3) in Appendix C.2 through the integration of $\mathbf{m}_{\mathrm{S}}$ into classical MU methods (FT, RL, GA, and IU).
@@
Extended from Fig. 4, we quantify the unlearning performance of Retrain, ESD, and SalUn using the two metrics introduced earlier, FID and UA, in Appendix C.3. With a comparable UA performance, it is worth highlighting that SalUn significantly outperforms ESD in FID. ESD seems to exhibit instability when forgetting and learning low-quality images like CIFAR-10. We also investigate the selection of the sparsity ratio for weight saliency in Appendix C.1.
@@ 8 Acknowledgement
C. Fan, J. Liu, and S. Liu were supported by the Cisco Research Faculty Award and the National Science Foundation (NSF) Robust Intelligence (RI) Core Program Award IIS-2207052. We would like to thank Jinghan Jia for the insightful discussions.
@@
Appendix A Pseudo code of RL-based SalUn.
Algorithm 1 Pseudo code of RL-based SalUn in classification tasks.
Hyper-parameters: learning rate $\eta$, mask threshold $\gamma$, and number of epochs $E$.
Relabeled forgetting set $\mathcal{D_{\mathrm{f}}}^{\prime}=\{(\mathbf{x}_{i},y^{\prime}_{i})|(\mathbf{x}_{i},y_{i})\in\mathcal{D_{\mathrm{f}}},y^{\prime}_{i}\neq y_{i}\}$
${{\bm{\theta}}_{\mathrm{u}}}\leftarrow{{\bm{\theta}}_{\mathrm{o}}}$
$\mathcal{D}^{\prime}\leftarrow\mathcal{D_{\mathrm{f}}}^{\prime}\cup\mathcal{D_{\mathrm{r}}}$
$\mathbf{g}_{\mathrm{S}}=\nabla_{{\bm{\theta}}}\mathbb{E}_{(\mathbf{x},y)\sim\mathcal{D}_{\mathrm{f}}}[\ell_{\mathrm{CE}}({\bm{\theta}};\mathbf{x},y)]|_{{\bm{\theta}}={{\bm{\theta}}_{\mathrm{o}}}}$ $\triangleright$ GA-based weight saliency from (5)
$\mathbf{m}_{\mathrm{S}}\leftarrow\mathds{1}\left(\left|\mathbf{g}_{\mathrm{S}}\right|\geq\gamma\right)$ $\triangleright$ Saliency mask from (3)
for $epoch\leftarrow 0\ldots E-1$ do
  for $\mathbf{b}\leftarrow$ all batches of $\mathcal{D}^{\prime}$ do
   $\mathbf{g}\leftarrow\nabla_{\bm{\theta}}\ell_{\texttt{SalUn}}^{(1)}({\bm{\theta}};\mathbf{b})|_{{\bm{\theta}}={{\bm{\theta}}_{\mathrm{u}}}}$ $\triangleright$ Batch-wise loss for (6)
   ${{\bm{\theta}}_{\mathrm{u}}}\leftarrow{{\bm{\theta}}_{\mathrm{u}}}-\eta\left(\mathbf{m}_{\mathrm{S}}\odot\mathbf{g}\right)$ $\triangleright$ One step SGD
  end for
end for
return ${\bm{\theta}}_{\mathrm{u}}$}
Algorithm 2 Pseudo code of RL-based SalUn in generation tasks.
Hyper-parameters: learning rate $\eta$, mask threshold $\gamma$, and number of iterations $T$.
Relabeled forgetting set $\mathcal{D_{\mathrm{f}}}^{\prime}=\{(\mathbf{x}_{i},c^{\prime})|(\mathbf{x}_{i},c_{i})\in\mathcal{D_{\mathrm{f}}},c^{\prime}\neq c_{i}\}$
${{\bm{\theta}}_{\mathrm{u}}}\leftarrow{{\bm{\theta}}_{\mathrm{o}}}$
$\mathcal{D}^{\prime}\leftarrow\mathcal{D_{\mathrm{f}}}^{\prime}\cup\mathcal{D_{\mathrm{r}}}$
$\mathbf{g}_{\mathrm{S}}=\nabla_{{\bm{\theta}}}\ell_{\mathrm{MSE}}({\bm{\theta}};\mathcal{D}_{\mathrm{f}})|_{{\bm{\theta}}={{\bm{\theta}}_{\mathrm{o}}}}$ $\triangleright$ GA-based weight saliency from (5)
$\mathbf{m}_{\mathrm{S}}\leftarrow\mathds{1}\left(\left|\mathbf{g}_{\mathrm{S}}\right|\geq\gamma\right)$ $\triangleright$ Saliency mask from (3)
for $it\leftarrow 0\ldots T-1$ do
  Sampling batch $\mathbf{b}$ from $\mathcal{D}^{\prime}$
  $\mathbf{g}\leftarrow\nabla_{\bm{\theta}}\ell_{\texttt{SalUn}}^{(2)}({\bm{\theta}};\mathbf{b})|_{{\bm{\theta}}={{\bm{\theta}}_{\mathrm{u}}}}$ $\triangleright$ Batch-wise loss for (7)
  ${{\bm{\theta}}_{\mathrm{u}}}\leftarrow{{\bm{\theta}}_{\mathrm{u}}}-\eta\left(\mathbf{m}_{\mathrm{S}}\odot\mathbf{g}\right)$ $\triangleright$ One step SGD
end for
return ${\bm{\theta}}_{\mathrm{u}}$}
Appendix B Soft-thresholding SalUn